### v1

@@ IV-A Destructive Self-Interference
Recall that RISs reflect the signals transmitted for non-target UEs as well. These reflections can impose destructive interference on corresponding UEs. To showcase some examples, Fig. 4-a and 4-b illustrate scenarios in a network where the RIS scattering matrix is to be tuned for the target user $\mathrm{U}_{1}$ operating at frequency $f_{1}$, which in turn affects a non-target user $\mathrm{U}_{2}$ operating at frequency $f_{2}$. Depending on the channel model assumption, various harmful self-interference effects for the non-target UE can be considered. Generally, the reflected rays from each array element to the non-target UE are added to the signal received from the direct path forming a multi-path fading channel. If the LoS channel model is considered, the beam pattern shaped by the RIS and potentially directed toward the non-target UE might increase the interference. If the non-target frequency $f_{2}$ is close to $f_{1}$, the reflected signal experiences negligible beam squint (Fig. 4-a). Otherwise, the main lobe of the reflected signal $x^{\prime}_{2}$ points in a different direction from that of the target reflected signal (Fig. 4-b). In either case, when the non-target UE lies within the main lobe of the reflected signal as shown in Fig. 4-a-b, QoS degradation is likely due to the potential incoherent multi-path arrival of the signal rays. While the RIS can be tuned to ensure the constructive arrival of signals from the direct path and reflected path, building optimal QoS for target UEs, this does not hold for non-target UEs at the same time.
@@ IV-B CSI Perturbation and Inaccuracy of Precoding
The scattering matrix of an RIS is continuously and frequently updated based on the CSI update scheduling considering changes of BS-to-RIS and RIS-to-target-UE channels without considering the impact on non-target UEs. Noting that the CSI estimation scheduling of the non-target UEs operating under different network protocols and managed by separate network coordinators are not synchronized with that of the target UEs, the estimated channels of non-target UEs become outdated as soon as the RIS scattering matrix is reconfigured for target UEs. This desynchronization leads to a reduction in the channel coherence time, making it unpredictable and thereby degrading the signal quality and connection reliability experienced by non-target UEs. For multi-operator MIMO communications, where each BS is supposed to serve multiple MIMO links at each dedicated carrier frequency, as depicted in Fig. 4-c, this together with the destructive self-interference for non-target UEs results in increasing decoding error probability due to inaccurate and out-dated computed precoders.
@@ IV-C Destructive Cross-Interference
This occurs when the RIS disrupts the spatial domain orthogonality of channel access for the UEs operating at the same carrier frequency. Consider a communication scenario shown in Fig. 4-d where $\mathrm{U}_{1}$ is the target user of the RIS and the non-target UEs ($\mathrm{U}_{2}$ and $\mathrm{U}_{3}$) are spatially separated and served by another operator on the same carrier frequency using spatial division multiple access (SDMA). Although $\mathrm{U}_{2}$ and $\mathrm{U}_{3}$ are considered to be sufficiently separated in the angular domain with respect to the corresponding transmitter to ensure effective SDMA performance, the deployment of the RIS may cause the signal $x_{2}$ intended for $\mathrm{U}_{2}$ to be reflected toward $\mathrm{U}_{3}$. As a result, cochannel cross-interference for $\mathrm{U}_{3}$ becomes a more challenging scenario compared to the previously discussed self-interference scenario. Unlike self-interference, which may occasionally be non-destructive, the cross-interference caused by RIS reflection is consistently destructive.
Figure 4: Illustration of various self-interference and cross-interference scenarios for non-target UEs in a single-operator multi-carrier (SO-MC), and multi-operator multi-carrier (MO-MC) network.
@@ V Numerical Results
In this section, we provide a simulation setup to evaluate the detrimental consequence of RIS deployment in multi-operator networks and show how deploying RISs to benefit the UEs served by some operators can harm other UEs located in the vicinity of the RIS served by other operators on carrier frequencies within the BoI of the RIS. We have considered a MIMO network with two operators each having one BS with $M=20$ antennas. We consider that BS1 and BS2 are served by operator 1 and operator 2 at frequencies $f_{1}=2.5$ GHz and $f_{2}=2.75$ GHz respectively. A two-dimensional network scenario is investigated according to Fig. 5 where an $N$-element diagonal RIS is considered to be located on 4 test positions, and the RIS phase shifters are adjusted by tuning the scattering matrix to maximize the weighted sum power of target-UEs [14, 7]. The impedance matrix is obtained from the ideal diagonal model presented in Fig. 2. The channel model and RIS model parameter values, as well as the solution scheme for finding the scattering matrix are adopted from [14, 7]. Here, we assume that the precoders of non-target UEs are tuned independently without the involvement of RIS, as the RIS is owned and controlled by the first operator to benefit its users. The tunning of the scattering matrix for target UEs operating at $f_{1}$ results in a change in the channel matrix of non-target UEs operating at $f_{2}$, in a mechanism expressed in Section III-B, which leads to the performance degradation for non-target UEs due to the increase in decoding error probability. It is seen from the figure that the RIS-induced performance degradation is relatively high, especially when the number of RIS elements gets higher and the RIS gets closer to non-target UEs. For example, considering a 70-element RIS located at (60,10)m, the average sum spectral efficiency for non-target UEs is decreased by a factor of 48% from 15.4 b/s/Hz to 8 b/s/Hz corresponding to RIS-interference-free (no RIS) and RIS-induced interference scenarios.
@@
Figure 5: The upper diagram shows the simulation setup scenario considering two BSs from two different operators, where the target and non-target UEs are serviced by operators 1 and 2 at carrier frequencies $f_{1}$ and $f_{2}$ respectively. The lower plot illustrates RIS-induced performance degradation of non-target UEs.

### v2

@@ IV-A Destructive Self-Interference
Recall that RISs also reflect the signals transmitted for non-target UEs. These reflections can impose destructive interference on corresponding UEs. To showcase some examples, Fig. 4-a and 4-b illustrate scenarios in a network where the RIS scattering matrix is to be tuned for the target user $\mathrm{U}_{1}^{(1)}$ operating at frequency $f_{1}$, which in turn affects the non-target user $\mathrm{U}_{2}^{(1)}$ operating at frequency $f_{2}$. Depending on the channel model assumption, various harmful self-interference effects for the non-target UE can occur. Generally, the reflected rays from each RIS element toward the non-target UE combine with the signal received from the direct path, forming a multi-path fading channel. If the LoS channel model is considered, the beam pattern shaped by the RIS and potentially directed toward the non-target UE might increase the interference. If the non-target frequency $f_{2}$ is close to $f_{1}$, the reflected signal experiences negligible beam squint (Fig. 4-a). Otherwise, the main lobe of the reflected signal $x^{\prime}_{2}$ points in a different direction from that of the target reflected signal (Fig. 4-b). In either case, when the non-target UE lies within the main lobe of the reflected signal as shown in Fig. 4-a and Fig. 4-b, QoS degradation is likely to happen due to the potential incoherent multi-path arrival of the signal rays. While the RIS can be tuned to ensure the constructive arrival of signals from the direct path and reflected path, achieving the desired QoS for target UEs may not result in the desired QoS for the non-target UEs.
@@ IV-B CSI Perturbation and Inaccuracy of Precoding
The scattering matrix of the RIS is continuously and frequently updated based on the CSI update schedule, considering changes in both BS-to-RIS and RIS-to-target-UE channels. This update process, however, does not account for potential negative impacts on non-target UEs. The CSI estimation schedules of non-target UEs, which operate under different network protocols and are managed by separate network coordinators, are not synchronized with those of the target UEs. As a result, the estimated channels of non-target UEs become outdated as soon as the RIS scattering matrix is reconfigured for the target UEs. This desynchronization reduces the channel coherence time, making it unpredictable and degrading the signal quality and connection reliability experienced by non-target UEs. For multi-operator MIMO communications, where each BS is supposed to serve multiple MIMO links at each dedicated carrier frequency, as depicted in Fig. 4-c, this together with the destructive self-interference for non-target UEs results in increasing decoding error probability due to inaccurate and outdated computed precoders.
@@ IV-C Destructive Cross-Interference
This occurs when the RIS disrupts the spatial domain orthogonality of channel access for the UEs operating at the same carrier frequency. Consider a communication scenario shown in Fig. 4-d where $\mathrm{U}_{1}^{(1)}$ is the target user of the RIS and the non-target UEs $\mathrm{U}_{2}^{(2)}$ and $\mathrm{U}_{3}^{(2)}$ are spatially separated and served by another access point (AP) 2 on the same carrier frequency using spatial division multiple access (SDMA). Although $\mathrm{U}_{2}^{(2)}$ and $\mathrm{U}_{3}^{(2)}$ are considered to be sufficiently separated in the angular domain with respect to the corresponding transmitter to ensure effective SDMA performance, the deployment of the RIS may cause the signal $x_{2}$ intended for $\mathrm{U}_{2}^{(2)}$ to be reflected toward $\mathrm{U}_{3}^{(2)}$. As a result, cochannel cross-interference for $\mathrm{U}_{3}^{(2)}$ becomes a more challenging scenario compared to the previously discussed self-interference scenario. Unlike self-interference, which may occasionally be non-destructive, the cross-interference caused by RIS reflection is consistently destructive.
Figure 4: Illustration of various self-interference and cross-interference scenarios for non-target UEs in a single-operator multi-carrier (SO-MC), and multi-operator multi-carrier (MO-MC) network, where $\mathrm{U}_{i}^{(m)}$ denotes the $i$-th user of AP $m\in\{1,2\}$.
@@ V Numerical Results
Figure 5: The upper diagram shows the simulation setup scenario considering two BSs from two different operators, where the target and non-target UEs are serviced by operators 1 and 2 at carrier frequencies $f_{1}=2.5$ GHz and $f_{2}=2.75$ GHz respectively. The middle plot illustrates the performance measure of RIS target UEs communicating with BS1. The lower plot shows the RIS-induced performance degradation of non-target UEs communicating with BS2.
In this section, we provide a simulation setup to evaluate the detrimental consequence of RIS deployment in multi-operator networks and show how deploying RISs to benefit the UEs served by a given operator can harm other UEs located in the vicinity of the RIS served by other operators on carrier frequencies within the BoI of the RIS. We considered a MIMO network with two operators each having one BS with $M=20$ antennas. We consider that BS1 and BS2 are served by operator 1 and operator 2 at frequencies $f_{1}=2.5$ GHz and $f_{2}=2.75$ GHz, respectively, and an RIS is used to assist the target UEs of BS1 whose LoS links are blocked. A two-dimensional network scenario is investigated according to Fig. 5-a where an $N$-element diagonal RIS is considered to be located on 3 test positions. The RIS scattering matrix is obtained by solving the optimization problem corresponding to the maximum weighted sum power of target UEs, where a zero-forcing BS precoder vector is employed for the UEs of each operator [7]. The RIS impedance matrix is obtained from the diagonal model in Fig. 2, where $R=1\ \Omega$, $L=2.5\ \mathrm{nH}$, and $L^{\prime}=0.7\ \mathrm{nH}$ have been considered. We employ a Rician fading channel model with Rician factor $K=1$ and an additive white Gaussian noise with variance $-30$ dBm. Here, we assume that the precoders of the non-target UEs are tuned independently without the involvement of the RIS, as the RIS is owned and controlled by the first operator to benefit its users. The tuning of the scattering matrix for the target UEs operating at $f_{1}$ results in a change in the channel matrix of the non-target UEs operating at $f_{2}$, in a mechanism expressed in Section III-B, which leads to the performance degradation for the non-target UEs due to the increase in decoding error probability.
The double-edged impact of RIS is evidenced by comparing Figs. 5-b and 5-c, where, for all scenarios, increasing the number of RIS elements enhances the performance of the target UEs (Fig. 5-b), while simultaneously exacerbating the negative impacts on the non-target UEs (Fig. 5-c). It is seen from the figure that the RIS-induced performance degradation is relatively high, especially when the number of RIS elements gets higher and the RIS gets closer to the non-target UEs. For example, considering a 40-element RIS located at (50,10) m, the average sum spectral efficiency of the non-target UEs is decreased by a factor of 49% from 15.2 b/s/Hz to 7.7 b/s/Hz corresponding to RIS-interference-free (no RIS) and RIS-induced interference scenarios, respectively.
As illustrated in the circuit model in Fig. 2, each RIS element can be modeled as a bandpass filter, where the variation in the circuit parameter values affects the BoI of the filter and also influences the beam squint for non-target UEs, as elaborated in Section III. To investigate the impact of RIS model parameter variations on system performance, we selected inductance $L$ as one of the model parameters and analyzed the performance of both target and non-target UEs for different values of $L$, as shown in Fig. 6. It is seen that the performance variation for both user groups is not monotonic. Furthermore, while better performance for target UEs corresponds to worse performance for non-target UEs for any given model parameter value $L$, the worst and best performance measures for target and non-target UEs, respectively, do not necessarily coincide with the same values of $L$.
Figure 6: Performance evaluation of target and non-target UEs versus $L$, a system model parameter influencing BoI and beam squint. A 50-element RIS is considered.